Circular No. 8382

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1187 SCORPII = NOVA SCORPII 2004 No. 2 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V1187 Sco has been given to this nova (cf. IAUC 8380, 8381).

D. K. Lynch, R. W. Russell, and D. Kim, Aerospace Corporation; M. Sitko, University of Cincinnati, and and R. B. Perry, Langley Research Center, NASA, report 3- to 13-μm spectroscopy of V1187 Sco using the Infrared Telescope Facility (+ BASS; 3′′.4-diameter beam) on Aug. 5.35 and 6.25 UT. At the spectral resolving power (30–100), the spectrum showed a fairly smooth continuum between 3 and 10 μm that is well-fitted by a 1500-K blackbody on Aug. 5 and by a 1400-K blackbody on Aug. 6. Atop the continuum, between 3.5 and 5 μm, were a number of poorly resolved emission lines that are believed to include Brα and Humphreys lines. Beyond 8 μm, the spectrum showed structured excess emission. Huα at 12.3 μm was evident, and it doubled in strength overnight. Also present, but as yet unidentified, were features at 8.6, 10.6 ([S IV] ?) and 11.2 μm, each of which also increased in strength between the Aug. 5 and 6. Infrared magnitudes: Aug. 5, L = 2.9, M = 2.4, N [10.5 μm] = 1.7; Aug. 6, 3.2, 2.7, 1.7. The spectrum is consistent with early development of novae near maximum light where the F-type pseudo-photospheric continuum gives way to an emission-line spectrum that is dominated by H I recombination lines and other low-excitation species. The infrared excess longward of 10 μm may represent thermal emission from pre-existing dust, because dust formation near maximum light is never observed.

Magnitude estimates (visual unless otherwise noted); Aug. 5.274 UT, 9.8 (J. Bedient, Honolulu, HI); 5.840, V = 9.63 (D. Hanžl, Brno, Czech Republic, 0.09-m reflector + CCD; comparison star HIP 85569); 5.843, 9.8 (K. Hornoch, Lelekovice, Czech Republic); 6.465, 9.8 (L. Kiss, Sydney, Australia); 6.836, 9.9 (Hornoch).

SUPERNOVA 2004do IN NGC 6708 L. A. G. Monard reports his discovery of an apparent supernova on

unfiltered CCD images taken on Aug. 4.841 (mag ∼ 16.3 ± 0.1) and 7.772 UT (mag ∼ 15.4 ± 0.1) with a 0.30-m reflector near Pretoria, South Africa. SN 2004do is located at α = 18^{h}55^{m}33^{s}.72, δ = −53^{o}43′22′′.8, which is 16′′ west and 3′′ north of the nucleus of NGC 6708. The new object is not visible on a red Digitized Sky Survey image (limiting mag 20.5) or on an image taken on 2004 June 10.87 by Monard (limiting mag 18.5).

2004 August 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green